Perpetual Inventory Using FIFO

Beginning inventory, purchases, and sales for Item Zeta9 are as follows:

Oct. 1 Inventory 47 units @ $19
7 Sale 37 units
15 Purchase 32 units @ $22
24 Sale 16 units

Assuming a perpetual inventory system and using the first-in, first-out (FIFO) method, determine (a) the cost of goods sold on October 24 and (b) the inventory on October 31.

a. Cost of goods sold on October 24 $
b. Inventory on October 31
Answer #1

a) Cost of goods sold on Oct 24

Total 16 units

10 units at a price on 1st Oct as per FIFO=10*19=$190

6 units at a price on 15th Oct as per FIFO=6*22=$132

Total = $322

b) Inventory on October 31= Opening stock +Purchases-Sales

= 47+32-37-16

=26 Units (costs=26*22=$572)